OpenStudy (anonymous):

The area of a circle varies directly as the square of the diameter. The area of a circle whose diameter is 20 inches is 314 square inches. What is the area of a circle whose diameter is 5 inches?

OpenStudy (shamim):

diameter d=5inch

OpenStudy (anonymous):

\[a=\pi (d/2)^2\]

OpenStudy (shamim):

area of a circlearea of a circle=πd^2/4

OpenStudy (anonymous):

area is pi r squared not d r=radius d= diameter

OpenStudy (anonymous):

@nincompoop please help bro

OpenStudy (shamim):

r=d/2

OpenStudy (anonymous):

yes

OpenStudy (shamim):

A=πr^2=π(d/2)^2

OpenStudy (anonymous):

ok so can we solve the problem

OpenStudy (shamim):

ya

OpenStudy (anonymous):

help me solve it please?

OpenStudy (anonymous):

ok i ampretty sure this is just a 20/314=5/x

OpenStudy (anonymous):

1570=20x

OpenStudy (anonymous):

1570/20=20x/20= 78.5=x

OpenStudy (anonymous):

are there choices? i would go with 78.5

OpenStudy (anonymous):

no choices :(

OpenStudy (anonymous):

can you check one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i could be completely wrong btw its 5:00 am here

OpenStudy (anonymous):

tired

OpenStudy (anonymous):

@ShadowLegendX could you help this guy and me?

OpenStudy (anonymous):

Neglecting air resistance, the distance an object falls from rest varies directly with the square of the time of fall. If an object falls 16.1 feet during one second, how far will it fall in 2.5 seconds?

OpenStudy (anonymous):

same thing, you try it this time

OpenStudy (anonymous):

2.5x16.1???

OpenStudy (anonymous):

no, first stack the object you dont know(x) on top of the other part you know: so x/2.5

OpenStudy (anonymous):

and then add an = and stack the other one with the same unit on the same side you had with you rfirst

OpenStudy (anonymous):

16.1/1 = x/2.5

OpenStudy (anonymous):

x/2.5=16.1/1

OpenStudy (anonymous):

yes either way works good job i didnt see it

OpenStudy (anonymous):

same difference

OpenStudy (anonymous):

and then cross multiply

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

oh man I never learned cross multiplying i never liked it

OpenStudy (anonymous):

its easy, what is 16.1x2.5?

OpenStudy (anonymous):

and what is 1*x

OpenStudy (anonymous):

wait isn't that what i asked in the beginning??? "2.5x16.1???"

OpenStudy (anonymous):

yea but if you wanna learn it you gotta write it out

OpenStudy (anonymous):

40.25/1=40.25=x

OpenStudy (anonymous):

well this is the way i look at it 1 second = 16.1 feet 2.5 seconds = 16.1 feet times 2.5 (because 2.5 is 2.5 times as much as 1) Is that right?

OpenStudy (anonymous):

its right i gues it just may not work in some situations.im not one to judge about that cause i am a lil wierd with math i just create things in my head to make it easier, i got suspended in 5th grade for making my teacher look bad lol

OpenStudy (anonymous):

haha well thanks so the final answer should be 40.3 right? since my answer has to be rounded to the tenth place

OpenStudy (anonymous):

sure, just remeber i could be wrong....also is this algebra 1?

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

I'm a Junior